OpenStudy (anonymous):

f(x)=x, if x is even 2x, if x is odd what is the value of f(8) +f(9)

OpenStudy (anonymous):

Since 8 is even and 9 is odd: \[ f(8)=8\\ f(9)=2(9)=18\implies\\ f(8)+f(9)=8+18=26 \]

OpenStudy (anonymous):

f(8) = 8 and f(9) = 18 so f(8) + f(9) = 8 + 18 = 26